Solve the recurrence relation using a recursion tree AND substitution method:

T(n) = T(n-1) + 10n
Answer #1
Recursion tree:
-------------------

Time complexity 
= T(0) + 10(1) + .... + 10(n-2) + 10(n-1) + 10n
 = 1+ 10(1+2+3+....+n) = 1+ 10(n(n+1)/2) = 1 + 5(n(n+1)) = O(n^2)

=================================

Substitution method:
------------------------
T(n) = T(n-1) + 10n
     = T(n-2) + 10(n-1) + 10n
     = T(n-3) + 10(n-2) + 10(n-1) + 10n
     ......
     ......
     ......
     = T(n-n) + 10(1) + .... + 10(n-2) + 10(n-1) + 10n
     = 1+ 10(1+2+3+....+n)
     = 1+ 10(n(n+1)/2)
     = 1 + 5(n(n+1))
     = O(n^2)